# Barycentric Neural Networks and Length-Weighted Persistent Entropy Loss: A Green Geometric and Topological Framework for Function Approximation

### v2

@@ 4 Barycentric Neural Networks (BNNs)
In this section, we test the capabilities of our proposed framework under the context of Resource-Efficient AI, assuming that we have limited resources for network design and training, such as a small number of base points and few training epochs. Our approach combine $\operatorname{BNN}$s with a new topological loss function, the length-weighted persistent entropy-based loss function, which, as its name suggests, is based on the persistent entropy variant proposed in Section 3, to approximate an objective function $f\colon[A,B]\to\mathbb{R}$, which is only known on a finite number of samples $S$, where $\{A,B\}\subset S\subset[A,B]$. A $\operatorname{BNN}$ is constructed from the limited set of base points, which define both its overall structure and its internal parameters. Consequently, $\operatorname{BNN}$s are non-trainable in the conventional sense, meaning that training in our framework amounts to iteratively adjusting the base points, thereby shaping the architecture. To guide this optimisation, we propose using a topological loss based on our novel variant of persistent entropy, the length-weighted persistent entropy ($\operatorname{LWPE}$), to guide the adjustment of base points and, consequently, the architecture of the $\operatorname{BNN}$s, optimising their distribution and ensuring effective approximation. We present experiments that show how our topological loss can be used to optimize the distribution of the base points that define $\operatorname{BNN}$s, approximating $f$ better and faster compared to using classical loss function, like mean squared error ($\operatorname{MSE}$), root mean squared error ($\operatorname{RMSE}$), mean absolute error ($\operatorname{MAE}$), or the LogCosh. We used the Tensorflow [^1] The TensorFlow documentation is available at https://www.tensorflow.org/. and Gudhi [^2] The Gudhi documentation is available at https://gudhi.inria.fr/. libraries to implement our proposed $\operatorname{BNN}$ and our proposed topological loss, as well as to conduct our experiments.
@@
In more realistic scenarios, the resources available for network design and training are often limited. For instance, it may be unrealistic to employ as many as 150 base points. Instead, we consider a function $f:[A,B]\subset\mathbb{R}\to\mathbb{R}$ that is represented by a point cloud, defined as a set $X=\{(x_{i},y_{i})\}_{i\in I}$, where each pair $(x_{i},y_{i})$ is a sample point, that is, $y_{i}=f(x_{i})$. In general, we say that a set $X$ is a point-cloud-based function if $x_{i}=x_{j}$ implies that $y_{i}=y_{j}$ for all $i,j\in I$. We aim to approximate $X$ using a $\operatorname{BNN}$ built from only a small, fixed number of base points, including the endpoints $A$ and $B$.
@@
Figure 4: Comparison of function approximation using a $\operatorname{BNN}$ with 8 base points, optimized via two persistent entropy-based loss functions. (a) Initial state with randomly initialized base points. (b) Final iteration (50) using standard persistent entropy-based loss function ($L_{\operatorname{PE}}$). (c) Final iteration using our length-weighted persistent entropy-based loss function ($L_{\operatorname{LWPE}}$). (d) MSE learning curves showing convergence differences between $L_{\operatorname{PE}}$ and $L_{\operatorname{LWPE}}$.
@@
Initially, we considered a topological loss function based on the standard persistent entropy:
$L_{\operatorname{PE}}=|\operatorname{PE}_{\operatorname{ref}}-\operatorname{PE}_{\operatorname{pred}}|$ (3)
@@
where both $\operatorname{PE}_{\operatorname{ref}}$ and $\operatorname{PE}_{\operatorname{pred}}$ are the persistent entropy computed via the lower-star filtrations of the reference and predicted point clouds, respectively. The objective of this persistent entropy-based loss function is to minimize the absolute difference between the persistent entropy derived from the reference and predicted point cloud, ensuring a similar topological structure. However, this standard persistent entropy-based approach presents a fundamental limitation, as discussed in Section 3, since it does not necessarily guide the $\operatorname{BNN}$ and its base points towards an accurate pointwise approximation of the target function. This limitation is clearly shown in Fig 4, where, after optimizing the base points to approximate $f(x)=\sin(x)$ on the interval [10,10][-10,10] using $L_{\mathrm{PE}}$, the $\operatorname{BNN}$ converges to a representation that preserves the relative proportions of the peaks but fails to align with their true amplitudes and oscillatory structure. Consequently, even though the persistent entropy values of the reference and prediction remain close, the learned function does not capture the actual maxima and minima of the target.
@@
To overcome this limitation, we employ our length-weighted persistent entropy ($\operatorname{LWPE}$), introduced in Section 3, as the basis for the topological loss function, replacing the standard persistent entropy. As shown in Fig. 4, optimizing base points using $\operatorname{LWPE}$ for computing the loss function instead of $\operatorname{PE}$ yields a significantly closer fit to the target function. This observation is further supported by the learning curves in Fig.4(d): while the standard entropy variant results to slower and limited convergence, our length-weighted version achieves rapid and stable minimization of the MSE, indicating more effective learning dynamics. Both experiments used SGD with a learning rate of 0.1 and limited resources: 50 epochs and 8 base points.
@@
$L_{\operatorname{LWPE}}=|\operatorname{LWPE}_{\operatorname{ref}}-\operatorname{LWPE}_{\operatorname{pred}}|.$ (4)
@@
To test our framework’s robustness, we evaluate its performance in approximating a noisy point cloud function using a limited and fixed number of base points. This scenario reflects more realistic applications, as in most real-world cases, point clouds will not be as smooth as the one presented previously. A key constraint in this setup is that, with a fixed number of points $n$, we can at most approximate a function (sampled by a point cloud $X$) whose persistence barcode contains at most $n/2$ bars, or equivalently, the $n/2$ most significant bars, if we assume that the function contains noise. As an illustrative example, refer to Fig. 5, where we consider a point cloud $X$ with 250 points representing a partial description of the sine function with noise (left). As discussed above, and by analyzing the obtained barcode derived from $X$, we can deduce that with $n=8$ points, we can at most approximate a function whose persistence barcode $D$ contains $n/2=4$ bars, or said in other words, we can approximate a barcode with the $n/2=4$ most longest bars of $D$ (considering the rest of the bars of $D$ as noise, as is represented in the right plot of Fig. 5).
Figure 6: Approximation of a noisy point-cloud-based function using a $\operatorname{BNN}$ with 8 base points, optimized via different loss functions. (a) Initial state with randomly placed base points. (b–f) Final epoch (50) BNN approximation using the following losses: (b) $\operatorname{MSE}$, (c) $\operatorname{RMSE}$, (d) $\operatorname{MAE}$, (e) LogCosh, (f) $L_{\operatorname{LWPE}}$.
@@
In Fig. 6 we can see an example where we aim to approximate a noisy point cloud function, utilizing $\operatorname{BNN}$s and our topological loss function $L_{\operatorname{LWPE}}$, as well as with more commonly used loss functions, such as $\operatorname{MSE}$, $\operatorname{RMSE}$, $\operatorname{MAE}$, or the LogCosh, under limited resources (8 base points and 50 epochs). With $L_{\operatorname{LWPE}}$, the $\operatorname{BNN}$ achieves a more effective distribution of base points and, consequently, a superior approximation under the same constraints. Furthermore, the learning curves in Fig. 7 show that optimizing the base points with our topological loss function $L_{\operatorname{LWPE}}$ exhibits a superior and faster performance than using classical loss functions. This indicates that $L_{\operatorname{LWPE}}$ provides more informative gradients during the early training stages, enabling a quicker and more accurate optimization of base point distributions, and consequently, a better approximation, reinforcing our framework’s effectiveness in handling complex and noisy data, even under limited settings.
@@
To showcase the applicability of our method to real-world data, we approximate a financial time series: the daily price of a Gold ETF during the year 2023 (point-cloud-based function consisting of 365 points), obtained using the yfinance python library [^3] The yfinance documentation is available at https://ranaroussi.github.io/yfinance/. Having a limited number of 30 base points, we aim to capture the signal’s essential structure, effectively treating this as a signal compression task that removes short-term noise. In Fig. 8, we observe that $L_{\operatorname{LWPE}}$ again outperforms the classical losses, reducing approximation error across all metrics within the first epochs, highlighting its suitability for approximation tasks and time-series compression with a limited number of base points and epochs.
In summary, using the $L_{\operatorname{LWPE}}$ to optimize the base points that define $\operatorname{BNN}$s demonstrates clear benefits in approximating complex, noisy functions with very few base points and epochs. This makes it especially appealing for real-world applications having limited resources.
@@ 6 Conclusions
In this work, we introduced a novel framework for function approximation that integrates geometric modeling with topological optimization. Central to our approach is the Barycentric Neural Network ($\operatorname{BNN}$), a shallow neural architecture whose structure and parameters are entirely determined by a fixed set of base points and their barycentric coordinates. This design enables exact representation of Continuous Piecewise Linear Functions ($\operatorname{CPLF}$s) and allows the $\operatorname{BNN}$ to serve as a flexible, interpretable, and structurally predefined architecture (model) for approximating general continuous functions. Unlike traditional neural networks, which require optimization of internal weights through backpropagation, our $\operatorname{BNN}$ shifts the learning process to the optimization of the base point positions themselves. To guide this optimization, we proposed a novel loss function based on the length-weighted persistent entropy ($\operatorname{LWPE}$). Our experimental results demonstrate that this $\operatorname{BNN}$-$\operatorname{LWPE}$ framework achieves superior performance for function approximation compared to using classical loss functions (MSE, RMSE, MAE, and log-cosh), especially under resource-constrained settings, such as a limited and fixed number of points, and limited training epochs, making it a strong candidate for applications aligned with the principles of Green AI. Our method offers a mathematically principled and computationally efficient alternative to standard deep learning approaches for function approximation, contributing to the development of more interpretable, sustainable, and theoretically grounded machine learning models.
Limitations
Despite its promising results, the proposed framework presents several limitations that merit discussion. A key challenge lies in the selection and initialization of the base points, which define the geometry of the $\operatorname{BNN}$. Because the network’s expressive power depends on these base points, poor initialization or insufficient point density can hinder convergence and compromise the quality of the approximation. The effectiveness of our topological loss is especially sensitive to the number and distribution of these points, which must be sufficient to capture the relevant information of the target function. Another limitation is linked to the nature of the function being approximated. Our topological loss is well-suited for functions with prominent and persistent topological structures, such as clear extrema or oscillatory behavior, but may underperform when applied to functions that are either highly irregular, noisy, or lack significant topological contrast. In such cases, the length-weighted persistent entropy-based loss function may fail to provide meaningful gradients, and statistical or hybrid loss functions might offer better alternatives. These limitations open avenues for future work on adaptive base point initialization strategies and hybrid loss functions that can combine topological and statistical information for more robust performance across a wider class of target functions.
@@
Future research focuses on extending this approach to higher-dimensional function approximation and exploring its applications in more real-world scenarios. In order to obtain the simplicial complex $\mathcal{K}$ that covers a compact set $\mathcal{C}\subset\mathbb{R}^{d}$, we could, for example, compute Delaunay triangulations [11] with a set of vertices $V$ being a set of points of $\mathbb{R}^{d}$. Although the time complexity of such construction is roughly $O(n^{\lfloor d/2\rfloor})$ (see [28]), it remains computationally feasible when $V$ is small enough. In addition, we aim to investigate adaptive initialization schemes for the base points, possibly incorporating prior information or combining topological and statistical heuristics. This would improve robustness and reduce the reliance on careful manual initialization. Furthermore, efforts will be directed toward formally proving the advantages of using our topological loss function. Additionally, the development of hybrid loss functions, which combine topological descriptors with classical metrics, may further enhance optimization in settings with noisy or weakly structured data. Another crucial direction is developing strategies to handle outliers in point-cloud-based function approximation, ensuring robustness against noisy data and improving generalization when dealing with real-world datasets.

### v3

@@ 4 Barycentric Neural Networks
In this section, we test the capabilities of our proposed framework ($\operatorname{BNN}$ trained with $\operatorname{LWPE}$-based loss) in the context of Resource-Efficient AI, assuming that we have limited resources for network design and training, such as a small number of base points and few training epochs. The goal is to approximate an objective function $f\colon[A,B]\to\mathbb{R}$, which is known on a finite set $S$ of points, where $\{A,B\}\subset S\subset[A,B]$, using just a limited number of base points to define the $\operatorname{BNN}$ that will approximate $f$.
In our approach, the optimization is entirely carried out on these base points. Once their positions are fixed, both the network’s connectivity and its associated parameters are uniquely determined. Each base point acts as a generator in the continuous input domain, not constrained to coincide with the discrete set $S$ where $f$ is known. To guide this optimization, we propose using a topological loss function based on the length-weighted persistent entropy ($\operatorname{LWPE}$). We compared it against classical losses, like $\operatorname{MSE}$, $\operatorname{RMSE}$, $\operatorname{MAE}$, or LogCosh, when used to optimize the same set of base points.
That is to say, the training follows a standard iterative procedure organized in epochs, but the variables being optimized are the positions of the base points. At epoch zero, an initial configuration of base points is generated, and the corresponding $\operatorname{BNN}$ is automatically constructed. Using this configuration, the network produces an initial approximation of the target function at the sampled points $S$. The error between the predicted and true values is then evaluated, either through classical losses such as $\operatorname{MSE}$, $\operatorname{RMSE}$, $\operatorname{MAE}$, or LogCosh, or through our proposed topological loss based on the $\operatorname{LWPE}$. For the topological loss, persistent homology is computed using the lower star filtration for both the target function and the $\operatorname{BNN}$ approximation. The corresponding persistence barcodes are then used to calculate their length-weighted persistent entropies. The difference between these two entropy values quantifies how well the approximation preserves the topological features of the target function. Once the loss is computed, the framework estimates how small displacements of the base points affect the overall error, updating their positions accordingly.
In summary, although no explicit weights are directly optimized, the movement of the base points within the continuous domain implicitly determines the configuration of the $\operatorname{BNN}$ parameters. This makes the training conceptually equivalent to that of traditional neural networks, but constrained to a compact and geometrically structured search space.
In summary, although no explicit weights are ever optimized, the movement of the base points within the continuous input domain implicitly determines the configuration of the $\operatorname{BNN}$ parameters. This makes the training process conceptually equivalent to the traditional optimization of neural networks, but confined to a compact and geometrically structured search space.
We used the TensorFlow [^1] The TensorFlow documentation is available at https://www.tensorflow.org/. and Gudhi [^2] The Gudhi documentation is available at https://gudhi.inria.fr/. libraries to implement our proposed $\operatorname{BNN}$ and our proposed topological loss, as well as to conduct our experiments.
@@
In more realistic scenarios, the resources available for network design and training are often limited. Instead, we consider a function $f:[A,B]\subset\mathbb{R}\to\mathbb{R}$ that is represented by a point cloud, defined as a set $X=\{(x_{i},y_{i})\}_{i\in I}$, where each pair $(x_{i},y_{i})$ is a sample point, that is, $y_{i}=f(x_{i})$. In general, we say that a set $X$ is a point-cloud-based function if $x_{i}=x_{j}$ implies that $y_{i}=y_{j}$ for all $i,j\in I$. We aim to approximate $X$ using a $\operatorname{BNN}$ built from a small set $S$ of base points that includes the endpoints $A$ and $B$.
@@
Figure 4: Comparison of function approximation using a $\operatorname{BNN}$ with 8 base points, optimized via two persistent entropy-based loss functions. (a) Initial state with randomly initialized base points. (b) 50th iteration, obtained using the standard persistent entropy-based loss $L_{\operatorname{PE}}$. (c) 50th iteration, obtained using the length-weighted persistent entropy-based loss $L_{\operatorname{LWPE}}$. (d) MSE learning curves showing convergence differences between $L_{\operatorname{PE}}$ and $L_{\operatorname{LWPE}}$.
5.1 $L_{\operatorname{PE}}$ vs $L_{\operatorname{LWPE}}$ topological loss
@@
Initially, we considered a topological loss based on the standard persistent entropy:
$L_{\operatorname{PE}}=|\operatorname{PE}_{\operatorname{ref}}-\operatorname{PE}_{\operatorname{pred}}|$
@@
where both $\operatorname{PE}_{\operatorname{ref}}$ and $\operatorname{PE}_{\operatorname{pred}}$ are the persistent entropy computed via the lower-star filtrations of the reference and predicted point clouds, respectively. The objective of this persistent entropy-based loss is to minimize the absolute difference between the persistent entropy derived from the reference and predicted point cloud, ensuring a similar topological structure. However, this standard persistent entropy-based approach presents a fundamental limitation, as discussed in Section 3, since it does not necessarily guide the $\operatorname{BNN}$ and its base points towards an accurate pointwise approximation of the target function. This limitation is clearly shown in Fig 4, where, after optimizing the base points to approximate $f(x)=\sin(x)$ on the interval [10,10][-10,10] using $L_{\mathrm{PE}}$, the $\operatorname{BNN}$ converges to a representation that preserves the relative proportions of the peaks but fails to align with their true amplitudes and oscillatory structure. Consequently, even though the persistent entropy values of the reference and prediction remain close, the learned function does not capture the actual maxima and minima of the target.
@@
To overcome this limitation, we employ our length-weighted persistent entropy ($\operatorname{LWPE}$), introduced in Section 3, as the basis for the topological loss, replacing the standard persistent entropy. As shown in Fig. 4, optimizing base points using $\operatorname{LWPE}$ loss instead of $\operatorname{PE}$ yields a significantly closer fit to the target function. This observation is further supported by the learning curves in Fig.4(d): while the standard entropy variant results to slower and limited convergence, our length-weighted version achieves rapid and stable minimization of the MSE, indicating more effective learning dynamics. Both experiments used the stochastic gradient descent (SGD) optimization algorithm with a learning rate of 0.1, as well as limited resources: 50 epochs and 8 base points.
@@
$L_{\operatorname{LWPE}}=|\operatorname{LWPE}_{\operatorname{ref}}-\operatorname{LWPE}_{\operatorname{pred}}|.$
@@
Figure 6: Approximation of a noisy point-cloud-based function using a $\operatorname{BNN}$ with 8 base points, optimized via different loss functions. (a) Initial state with randomly placed base points. (b–f) Final epoch (50) BNN approximation using the following loss functions:(b) $\operatorname{MSE}$, (c) $\operatorname{RMSE}$, (d) $\operatorname{MAE}$, (e) LogCosh, (f) $L_{\operatorname{LWPE}}$.
5.2 Robustness under noise and a limited number of base points
To test the robustness of our framework, we evaluate its performance in approximating a noisy point cloud function using a limited number of base points. This scenario reflects more realistic applications, as in most real-world cases, point clouds will not be as smooth as the one presented previously.
First, as an illustrative example of Remark 5, see Fig. 5, where we consider a point cloud $X$ with 250 points representing a noisy partial description of the sine function (left). As discussed above, and by analyzing the obtained persistence barcode computed from $X$, we can deduce that with $n=8$ base points, we can approximate the sine function with the $n/2=4$ longest bars, while treating the remaining bars of the barcode as noise (see the right plot of Fig. 5).
Second, in Fig. 6, we can see an example where we aim to approximate a noisy version of the sine function represented by a noisy point cloud, using limited resources (8 base points and 50 epochs), utilizing BNNs. We compared the results obtained with the $L_{\operatorname{LWPE}}$ topological loss against more commonly used loss functions, such as $\operatorname{MSE}$, $\operatorname{RMSE}$, $\operatorname{MAE}$, or the LogCosh. We have observed that with $L_{\operatorname{LWPE}}$, the $\operatorname{BNN}$ achieves a more effective distribution of base points and, consequently, a superior approximation under the same constraints. Furthermore, the learning curves in Fig. 7 show that optimizing the base points with the $L_{\operatorname{LWPE}}$ loss exhibits a superior and faster performance than using classical loss functions. This indicates that the $L_{\operatorname{LWPE}}$ loss provides more informative gradients during the early training stages, enabling a quicker and more accurate optimization of base point distributions, and consequently, a better approximation, reinforcing our framework’s effectiveness in handling complex and noisy data, even under limited settings.
@@ 5.2 Robustness under noise and a limited number of base points
5.3 Robustness under outliers
A further challenge in real-world data is the presence of outliers, i.e., extreme values that deviate significantly from the underlying structure of the data. To test the robustness of our framework under this setting, we consider again a noisy sine function (similar to Subsection 5.2), but we introduce some artificial outliers at some of its peaks. This creates a distorted point cloud where the global shape of the sine curve coexists with an anomalous observation that could potentially bias the learning process.
When trained exclusively with our topological loss based on $\operatorname{LWPE}$, the model tends to overfit to the outlier. This behavior is expected due to the sensitivity of $\operatorname{LWPE}$ to outliers: the outlier generates a highly persistent topological feature (e.g., a long bar in the persistence diagram), which dominates the $\operatorname{LWPE}$ computation and biases the optimization toward reproducing the anomalous value rather than the true underlying signal.
To address this limitation, we propose a hybrid training strategy that combines $\operatorname{LWPE}$ as a regularizer with a classical loss such as MSE. As shown in Fig. 8, this composite objective effectively balances local fidelity with global structural consistency. The result avoids two pitfalls: (1) the slow convergence and poor shape recovery observed when using MSE alone (as shown in Section 5.2), and (2) the outlier-driven distortion caused by relying solely on $\operatorname{LWPE}$. The hybrid approach thus achieves a high-quality approximation that respects both the fine-grained data and the overarching topology of the target function.
Figure 8: Approximation of a noisy point-cloud-based function containing outliers using a $\operatorname{BNN}$ with 8 base points, optimized using $L_{\operatorname{LWPE}}$ as a regularizer of MSE. On the left, the initial state with randomly placed base points, and on the right, the final BNN approximation after 50 epochs.
These findings underscore the robust performance even in the presence of data outliers, a common reality in practical applications. Moreover, the modular design of our framework allows the topological loss to be flexibly incorporated as a regularizer, enabling practitioners to tailor the trade-off between performance and structural integrity according to the demands of their specific domain.
5.4 Testing our method with real-world data
To showcase the applicability of our method to real-world data, we approximate a financial time series: the daily price of a Gold ETF during the year 2023 (point-cloud-based function consisting of 365 points), obtained using the yfinance Python library [^3] The yfinance documentation is available at https://ranaroussi.github.io/yfinance/. Having a limited number of 30 base points, we aim to capture the signal’s essential structure, effectively treating it as a signal compression task that removes short-term noise. In Fig. 9, we observe that $L_{\operatorname{LWPE}}$ again outperforms the classical loss functions, reducing approximation error across all metrics within the first epochs, highlighting its suitability for approximation tasks and time-series compression with a limited number of base points and epochs.
@@ 5.4 Testing our method with real-world data
In summary, using the $L_{\operatorname{LWPE}}$ loss to optimize the base points that define $\operatorname{BNN}$s demonstrates clear benefits in approximating complex, noisy functions with very few base points and epochs. This makes the approach particularly attractive for real-world applications with limited computational resources.
@@ 6 Conclusions
In this work, we introduced a novel framework for function approximation that integrates geometric modeling with topological optimization. At the core of our approach lies the Barycentric Neural Network ($\operatorname{BNN}$), a shallow neural architecture whose structure and parameters are entirely determined by a set of trainable base points and their barycentric coordinates. This design enables exact representation of Continuous Piecewise Linear Functions ($\operatorname{CPLF}$s), and therefore, $\operatorname{BNN}$s serve as a flexible, interpretable, and structurally predefined architecture for approximating general continuous functions. Unlike traditional neural networks, which require optimization of internal weights, our $\operatorname{BNN}$ shifts the learning process to the optimization of the base points. To guide this optimization, we proposed a novel topological loss function based on the length-weighted persistent entropy. Our experimental results demonstrate that this $\operatorname{BNN}$-$\operatorname{LWPE}$ framework achieves superior performance for function approximation, including noisy functions, compared to using classical loss functions (MSE, RMSE, MAE, and LogCosh), especially under resource-constrained settings, such as a limited and fixed number of base points, and limited training epochs, making it a strong candidate for function approximation, as well as following with the principles of Green AI. In summary, our method provides a mathematically principled and computationally efficient alternative to standard deep learning approaches for function approximation, contributing to the development of more interpretable, sustainable, and theoretically grounded machine learning models.
@@
Future research focuses on extending this approach to higher-dimensional function approximation and exploring its applications in more real-world scenarios. To obtain the simplicial complex $\mathcal{K}$ that covers a compact set $\mathcal{C}\subset\mathbb{R}^{d}$, we could, for example, compute Delaunay triangulations [11] with a set of vertices $V$ being a set of points of $\mathbb{R}^{d}$. Although the time complexity of such construction is roughly $O(n^{\lfloor d/2\rfloor})$, it remains computationally feasible when $V$ is small enough (see [29]). In addition, we aim to investigate adaptive initialization schemes for the base points, possibly incorporating prior information or combining topological and statistical heuristics. Such approaches could enhance robustness and reduce dependence on careful manual initialization. Notably, our experiments already demonstrate that, whether base points are initialized uniformly or randomly, our topological loss consistently yields a better optimization process compared to classical losses. Furthermore, efforts will be directed toward formally proving the advantages of using the topological loss $\operatorname{LWPE}$. Additionally, the development of hybrid loss functions, which combine topological descriptors with classical metrics, may further enhance optimization in settings with noisy or weakly structured data, or containing outliers, as we checked in Subsection 5.3 where we show how our method can be adapted to be robust to outliers. Finally, while the current BNN architecture is designed to exactly represent CPLFs, future work will explore extending this framework to approximate smoother functions by allowing nonlinear interpolations between base points. This generalization would enable the model to capture higher-order smoothness while maintaining its geometric interpretability and parameter efficiency. Moreover, such an approach could naturally generalize to multivariate settings, broadening the applicability of the proposed method.